# August 6, 2024

### v1

@@ Abstract
The real part of pseudo entropy in conformal field theories is holographically calculated by the area of some extremal spacelike surfaces in the dual dS and AdS spacetimes. We show that the flat-space limit of these curves in three dimensional (A)dS spacetimes is well-defined. We find that if the length of the curves is calculated from the radial coordinate where the retarded time is extremum, then after taking the flat-space limit, the entanglement entropy of the dual theory of three-dimensional flat spacetime is obtained. For dS spacetime, the radial coordinate corresponding to the extremum of retarded time is located inside the cosmlogical horizon. Our results suggest that on the field theory side, the entanglement entropy in the dual theory of flat spacetimes should be obtained from the ultra-relativistic limit of pseudo entropy in the dual CFT to (A)ds spacetimes.
@@ 1 Introduction
It has been proposed in [1]-[2] that the flat-space limit (i.e., the zero cosmological constant limit) of asymptotically AdS spacetimes is equivalent to the ultra-relativistic limit [3] in conformal field theory. As a result, it has been suggested that the dual field theory to asymptotically flat spacetimes in three dimensions is a two-dimensional ultra-relativistic field theory, known as Carrollian Conformal Field Theory (CCFT) [2, 4, 5]. This theory exists at null infinity of asymptotically flat spacetimes, and this duality is referred to as the flat/CCFT correspondence.  [^1] In some papers, this duality is also called flat/BMSFT. BMSFT refers to a field theory with BMS symmetry, which is the asymptotic symmetry of asymptotically flat spacetimes at null infinity [6]-[9]. This symmetry is equivalent to Carrollian conformal symmetry in one lower dimension.
@@
Typically, a general system is decomposed into subsystems at constant time levels. This allows the Hilbert space to be expressed as a tensor product of its subspaces. From a general density matrix in the original system, one can derive a reduced density matrix for spacelike subsystems and calculate the entanglement entropy. For CFTs, the holographic dual of this entanglement entropy, according to Ryu-Takayanagi [10], is represented by extremal spacelike surfaces in the bulk of asymptotically AdS spacetimes. Recently, [11]-[12] have studied the generalization of entanglement entropy for timelike subsystems and proposed holographic duals for these as well. This entropy, known as timelike entanglement entropy, is a complex quantity where the real part corresponds to the area of spacelike extremal surfaces and the imaginary part corresponds to the area of timelike extremal surfaces in the bulk. In asymptotically AdS spacetimes, these spacelike surfaces are divided into several parts connected to the subsystem at the boundary but not directly to each other in the bulk, with timelike surfaces connecting them. Furthermore, [11]-[12] argue that timelike entanglement entropy is essentially a special case of pseudo-entropy [13], where instead of the reduced density matrix, the transition matrix is used (see section 2.1).
The Ryu-Takayanagi proposal for holographic entanglement entropy can be explored in other dualities. For asymptotically de Sitter (dS) spacetimes, this correspondence is known as dS/CFT, suggesting that a (d-1)-dimensional Euclidean conformal field theory is the holographic dual of d-dimensional asymptotically dS spacetimes [14]. This field theory, defined at the timelike infinity of dS spacetime, is non-unitary, so the entanglement entropy for its subsystems is essentially pseudo-entropy and must be complex. The holographic dual of this pseudo-entropy is given in [11]-[12], consisting of extremal surfaces in the bulk, with parts being spacelike and others timelike. Unlike in AdS, the timelike surfaces here connect to the subsystem at timelike infinity, while the spacelike surfaces appear inside the bulk and connect to the timelike surfaces.
For CCFTs, as the dual theory of asymptotically flat spacetimes, entanglement entropy can also be defined. Using the infinite-dimensional symmetry of two-dimensional CCFTs, a universal formula for the entanglement entropy of intervals is provided in [15]. The holographic interpretation of this entropy is proposed in [16] as extremal curves in asymptotically flat spacetimes. These curves consist of spacelike and null parts, with the spacelike part entirely in the bulk and the null part connecting the ends of the interval at null infinity to the spacelike curve.
@@ 2.1 Entanglement entropy and pseudo entropy in field theory
If subsystem A is a timelike region of the total system, the reduced density matrix can still be defined for it using formula (2.1), and the entanglement entropy can be obtained using formula (2.2). However, in this case, the entropy is complex-valued. This entropy is reffered to as timelike entanglement entropy. The complexity arises because the density matrix in this case is not Hermitian. In the paper [12], it is argued that the timelike entanglement entropy can be reformulated using a transition matrix. Therefore, the timelike entanglement entropy itself is considered a pseudo entropy.
@@ 2.2 Entanglement entropy in the dual theory of flat spacetimes
In paper [1]-[2] , it is proposed that the dual field theory to asymptotically flat spacetime in three dimensions is a two-dimensional CCFT, which possesses the following infinite-dimensional symmetry:
@@
It is evident from algebra (2.2) that the symmetries of two-dimensional CCFTs are infinite dimensional, similar to those of two-dimensional CFTs . This infinite-dimensional symmetry allow for the identification of universal properties within these theories. Among these features are universal formulas for entanglement entropy. Research into this aspect of two-dimensional CCFTs began in [15] , where it it was shown that, in a CCFT on the plane, if we consider the subsystem to be the interval $\left(-\dfrac{l_{u}}{2}<u<\dfrac{l_{u}}{2},\,-\dfrac{l_{\phi}}{2}<\phi<\dfrac{l_{\phi}}{2}\right)$ , the entanglement entropy is given by
@@
where $\epsilon$ is a cutoff used to regulate the interval. For CCFTs on the cylinder this formula is given by
@@
The holographic interpretation of entanglement entropy was proposed in [16]. The algebra (2.2) on the gravity side corresponds to the asymptotic symmetry of asymptotically flat spacetimes at null infinity [8]. Consequently, dual CCFTs must be defined at null infinity, and each subsystem in these theories will correspond to intervals at null infinity as well. This implies that the extremal spacelike curves , whose length leads to the entropies (2.6) and (2.7) , do not connect the two ends of the interval directly. In fact, the proposal of [16] is that the holographic interpretation in this context involves spacelike curve whose endpoints are connected to the two ends of the interval at null infinity by two null rays.
@@ 3 Holographic pseudo entropy in AdS and its flat limit
In this section we aim to find the holographic timelike entanglement entropy for an interval on the boundary of AdS3 and then take the flat-space limit. As an example of pseudo entropy in CFT, this entropy is inherently complex-valued . The holographic dual of this entropy introduced in [11]-[12] consists of two geodesics: one spacelike and one timelike. To obtain the flay-space limit, we only need the spacelike part. We begin by considering a spacelike interval on the boundary of AdS3, compute its holographic entanglement entropy, and then perform a Wick rotation to extract the spacelike part of the holographic timelike entanglement entropy.
@@ 4 Holographic pseudo entropy in dS and its flat-space limit
To find extremal curves with a well-defined flat-space limit in dS spacetimes, we use the fact that the metric of dS3 can be obtained from the metric of AdS3 by replacing $\ell^{2}\to-\ell^{2}$. Thus, we start form dS3 written in the BMS coordinate as
@@ 5 Summary and future directions
Another issue discussed in this article is that to accurately obtain the entanglement entropy in CCFT, one should calculate only the length of a segment of the spacelike curve rather than its entirety. This segment length is calculated from a point in dS or AdS spacetime where the time direction, $u$, is extremum. The reason for this and its significance in dual field theories are questions we aim to answer in future studies. This problem becomes more intriguing when we consider formulas (3.23) and (4.11), which result from calculating the segment length of spacelike curves before taking the flat-space limit. These two formulas, which can be transformed into each other by the $\ell\to i\ell$ transformation, are quantitative expressions in Lorentzian CFT and Euclidean CFT, respectively, whose flat-space limit yields the entanglement entropy of CCFT. Understanding this quantity and its relation to pseudo-entropy or entanglement entropy is also a compelling question that will be the focus of our future researches.
Acknowledgments

### v2

# October 6, 2024

@@ Abstract
The real part of pseudo entropy in conformal field theories is holographically calculated by the area of some extremal spacelike surfaces in the dual dS and AdS spacetimes. We show that the flat-space limit of these curves in three-dimensional (A)dS spacetimes is well defined. We find that if the length of the curves is calculated from the radial coordinate where the retarded time is extremum, then after taking the flat-space limit, the entanglement entropy of the dual theory of three-dimensional flat spacetime is obtained. For dS spacetime, the radial coordinate corresponding to the extremum of retarded time is located inside the cosmological horizon. Our results suggest that on the field theory side, the entanglement entropy in the dual theory of flat spacetimes should be obtained from the ultrarelativistic limit of pseudo entropy in the dual CFT to (A)dS spacetimes.
@@ 1 Introduction
It has been proposed in [1]-[2] that the flat-space limit (i.e., the zero cosmological constant limit) of asymptotically AdS spacetimes is equivalent to the ultrarelativistic limit [3] in conformal field theory. As a result, it has been suggested that the dual field theory to asymptotically flat spacetimes in three dimensions is a two-dimensional ultra-relativistic field theory, known as Carrollian Conformal Field Theory (CCFT) [2, 4, 5]. This theory exists at null infinity of asymptotically flat spacetimes, and this duality is referred to as the flat/CCFT correspondence.  [^1] In some papers, this duality is also called flat/BMSFT. BMSFT refers to a field theory with BMS symmetry, which is the asymptotic symmetry of asymptotically flat spacetimes at null infinity [6]-[9]. This symmetry is equivalent to Carrollian conformal symmetry in one lower dimension.
@@
Typically, a general system is decomposed into subsystems at constant time levels. This allows the Hilbert space to be expressed as a tensor product of its subspaces. From a general density matrix in the original system, one can derive a reduced density matrix for spacelike subsystems and calculate the entanglement entropy. For CFTs, the holographic dual of this entanglement entropy, according to Ryu-Takayanagi [10], is represented by extremal spacelike surfaces in the bulk of asymptotically AdS spacetimes. Recently, the authors of [11]-[12] have studied the generalization of entanglement entropy for timelike subsystems and proposed holographic duals for these as well. This entropy, known as timelike entanglement entropy, is a complex quantity where the real part corresponds to the area of spacelike extremal surfaces and the imaginary part corresponds to the area of timelike extremal surfaces in the bulk. In asymptotically AdS spacetimes, these spacelike surfaces are divided into several parts connected to the subsystem at the boundary but not directly to each other in the bulk, with timelike surfaces connecting them. Furthermore, [11]-[12] argue that timelike entanglement entropy is essentially a special case of pseudo-entropy [13], where instead of the reduced density matrix, the transition matrix is used (see section 2.1).
The Ryu-Takayanagi proposal for holographic entanglement entropy can be explored in other dualities. For asymptotically de Sitter (dS) spacetimes, this correspondence is known as dS/CFT, suggesting that a (d-1)-dimensional Euclidean conformal field theory is the holographic dual of d-dimensional asymptotically dS spacetimes [14]. This field theory, defined at the timelike infinity of dS spacetime, is non-unitary, so the entanglement entropy for its subsystems is essentially pseudo-entropy and must be complex. The holographic dual of this pseudo-entropy is given in [11]-[12], consisting of extremal surfaces in the bulk, with parts being spacelike and others timelike (see also [15]-[19]). Unlike in AdS, the timelike surfaces here connect to the subsystem at timelike infinity, while the spacelike surfaces appear inside the bulk and connect to the timelike surfaces.
For CCFTs, as the dual theory of asymptotically flat spacetimes, entanglement entropy can also be defined. Using the infinite-dimensional symmetry of two-dimensional CCFTs, a universal formula for the entanglement entropy of intervals is provided in [20] (see also the recent work [21]). The holographic interpretation of this entropy is proposed in [22] as extremal curves in asymptotically flat spacetimes. These curves consist of spacelike and null parts, with the spacelike part entirely in the bulk and the null part connecting the ends of the interval at null infinity to the spacelike curve.
@@ 2.1 Entanglement entropy and pseudo entropy in field theory
If subsystem A is a timelike region of the total system, the reduced density matrix can still be defined for it using formula (2.1), and the entanglement entropy can be obtained using formula (2.2). However, in this case, the entropy is complex-valued. This entropy is referred to as timelike entanglement entropy. The complexity arises because the density matrix in this case is not Hermitian. In the paper [12], it is argued that the timelike entanglement entropy can be reformulated using a transition matrix. Therefore, the timelike entanglement entropy itself is considered a pseudo entropy.
@@ 2.2 Entanglement entropy in the dual theory of flat spacetimes
In papers [1]-[2] , it is proposed that the dual field theory to asymptotically flat spacetime in three dimensions is a two-dimensional CCFT, which possesses the following infinite-dimensional symmetry:
@@
It is evident from algebra (2.2) that the symmetries of two-dimensional CCFTs are infinite dimensional, similar to those of two-dimensional CFTs . This infinite-dimensional symmetry allows for the identification of universal properties within these theories. Among these features are universal formulas for entanglement entropy. Research into this aspect of two-dimensional CCFTs began in [20], where it was shown that, in a CCFT on the plane with coordinates $(u,\phi)$ where $u$ is known as retarded time [8] and $\phi$ is non-periodic, if we consider the sub-system to be a line with end points $\left(u=-\dfrac{l_{u}}{2},\phi=-\dfrac{l_{\phi}}{2}\right)$ and $\left(u=\dfrac{l_{u}}{2},\phi=\dfrac{l_{\phi}}{2}\right)$ where $l_{u}$ and $l_{\phi}$ are constants, the entanglement entropy is given by
@@
where $\epsilon$ is a cutoff used to regulate the interval and $c_{L}$ and $c_{M}$ are central charges of (2.2). For CCFTs on the cylinder with coordinates $(u,\phi)$ where $\phi$ is periodic, this formula is given by
@@
The holographic interpretation of entanglement entropy was proposed in [22]. The algebra (2.2) on the gravity side corresponds to the asymptotic symmetry of asymptotically flat spacetimes at null infinity [8]. Consequently, dual CCFTs must be defined at null infinity, and each subsystem in these theories will correspond to intervals at null infinity as well. This implies that the extremal spacelike curves , whose length leads to the entropies (2.6) and (2.7) , do not connect the two ends of the interval directly. In fact, the proposal of [22] is that the holographic interpretation in this context involves spacelike curves whose end points are connected to the two ends of the interval at null infinity by two null rays.
@@ 3 Holographic pseudo entropy in AdS and its flat limit
In this section we aim to find the holographic timelike entanglement entropy for an interval on the boundary of AdS3 and then take the flat-space limit. As an example of pseudo entropy in CFT, this entropy is inherently complex-valued . The holographic dual of this entropy introduced in [11]-[12] consists of two geodesics: one spacelike and one timelike. To obtain the flat-space limit, we only need the spacelike part. We begin by considering a spacelike interval on the boundary of AdS3, compute its holographic entanglement entropy, and then perform a Wick rotation to extract the spacelike part of the holographic timelike entanglement entropy.
@@ 4 Holographic pseudo entropy in dS and its flat-space limit
To find extremal curves with a well-defined flat-space limit in dS spacetimes, we use the fact that the metric of dS3 can be obtained from the metric of AdS3 by replacing $\ell^{2}\to-\ell^{2}$. Thus, we start from dS3 written in the BMS coordinate as
@@ 5 Summary and future directions
Another issue discussed in this article is that to accurately obtain the entanglement entropy in CCFT, one should calculate only the length of a segment of the spacelike curve rather than its entirety. This segment length is calculated from a point in dS or AdS spacetime where the time direction, $u$, is extremum. The reason for this and its significance in dual field theories are questions we aim to answer in future studies. This problem becomes more intriguing when we consider formulas (3.23) and (4.11), which result from calculating the segment length of spacelike curves before taking the flat-space limit. These two formulas, which can be transformed into each other by the $\ell\to i\ell$ transformation, are quantitative expressions in Lorentzian CFT and Euclidean CFT, respectively, whose flat-space limit yields the entanglement entropy of CCFT. Understanding this quantity and its relation to pseudo-entropy or entanglement entropy is also a compelling question that will be the focus of our future research.
Acknowledgment